justbored:

Which of the following is equivalent to - 4(2x - 3)? A. -12+8x B. 12+8x C. 12-8x D. 12x+8

Extrinix:

Do substitution, \((-4\times2x)\) \((-4\times-3)\) solve these and you will have your equation.

Extrinix:

Correct!

MunceeDaSteppa:

Right!

loverofmine:

-4*2 = -8 -4*-3 = 12 so, it is 12-8x :)